Even though most corporate bonds in the United States make coupon payments semiannually, bonds issued elsewhere often have annual coupon payments. Suppose a German company issues a bond with a par value of €1,000, 20 years to maturity, and a coupon rate of 6.2 percent paid annually. If the yield to maturity is 7.3 percent, what is the current price of the bond? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Price € This
Answer #1

Annual coupon=1000*6.2%=62

Hence current price=Annual coupon*Present value of annuity factor(7.3%,20)+1000*Present value of discounting factor(7.3%,20)

=62*10.35142214+1000*0.244346183

=886.13(Approx).

NOTE:

1.

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=62[1-(1.073)^-20]/0.073

=62*10.35142214

2.Present value of discounting factor=1000/1.073^20

=1000*0.244346183